# Kinetic Theory of Quasiparticles, Retarded Correlators and Hydrodynamics

### v1

@@ Abstract
Kinetic theory provides a powerful framework for investigating macroscopic and microscopic properties of weakly and strongly interacting plasmas. By assuming a constant mass profile within the kinetic theory approach, we examine the collective behavior of a system of massive relativistic particles within the Maxwell-Boltzmann equilibrium distribution. We derive the two-point retarded correlation functions for both charge and energy-momentum tensor components at arbitrary momentum and frequency values. To obtain analytical results, we expand relations for very small and large mass-to-temperature ratios ($m/T$). For large ratios, no significant effects are observed in the charge, shear, or sound propagation channels. However, for small ratios in the charge correlation function within the hydrodynamic limit, the mass tends to shift the hydrodynamic poles toward the unstable region, i.e., $\text{Im}(\omega)\geq 0$. This instability emerges at a critical value of $(m/T)_{c}\simeq 1.234$. Furthermore, in the sound channel, the mass amplifies the standard propagating sound mode, converting it into a purely (positive) imaginary mode. Consequently, all modes in this channel become purely imaginary, with instability becoming a dominant characteristic. In contrast, in the shear channel, all modes converge to their massless limits. Similar to the massless case, a threshold value exists for $(k\tau)$, below which the correlators exhibit solutions. This behavior originates from the presence of a logarithmic branch cut. For larger momenta, solutions appear significantly below the cut, giving rise to non-hydrodynamic modes. We find that hydrodynamic poles arise in the strong coupling regime, while the weak coupling regime features a logarithmic branch cut between $\omega=k$ and $\omega=-k$ that is analogous to the massless scenario. Additionally, we derive the shear and bulk viscosity transport coefficients, along with other coefficients up to third order in gradients, as an expansion in terms of small $(m/T)$.
@@ I Introduction
In this paper, we investigate the Boltzmann equation within the Relaxation Time Approximation (RTA) for a system of particles with a constant mass profile. The chemical potential of the particles is neglected in the background limit. Employing a variational approach, we derive retarded two-point correlation functions for the charge and energy-momentum tensor components at arbitrary frequencies and momenta. To explore collective behaviors, we examine the small-momentum limits of these functions. For the sake of analytical tractability, we expand expressions for both small and large values of the mass-to-temperature ratio, $x=m/T$. Our results demonstrate that for very large ratios, collective behavior is absent in the charge, shear, and sound channels. In contrast, small $x$ exhibits intriguing characteristics. Specifically, in the charge correlation functions, the mass tends to drive hydrodynamic modes toward the unstable region, characterized by $\mbox{Im}(\omega)\geq 0$, across small momentum values. At $x\sim 1.234$, the diffusive pole in the charge correlators begins to reside in the upper-half of the complex plane. This phenomenon arises because the mass alters the transport coefficients, thereby influencing the stability of the system near the thermal equilibrium state. In the energy-momentum sector, we derive the fluctuations of the energy-momentum components and the complete spectrum of retarded correlation functions. We demonstrate that, up to certain contact terms, the Ward identities are satisfied. Furthermore, we compute the shear and bulk transport coefficients as well as other ones up to the third order in derivatives for a non-conformal fluid. It has been shown by these computations that there is a critical $x$ beyond which the stability or the second law of thermodynamics is violated. The hydrodynamic poles in the shear and sound channels are identified, revealing that in the sound channel, all modes become purely imaginary and susceptible to instability. This indicates the absence of sound wave propagation in a massive plasma. Additionally, we observe the emergence of a branch cut between the $\omega=k$ and $\omega=-k$ in the solutions in the weak coupling regime ($\tau_{eq}T\to\infty$), while in the strong coupling regime ($\tau_{eq}T\to\infty$), hydrodynamic poles arises above a critical $(k\tau_{eq})$. This is the consequence of the logarithmic branch cut for small momenta. Although the structure and arrangement of the branch cut and poles resemble those in the massless case, the presence of mass shifts the hydrodynamic poles into the upper half of the complex plane.
@@ II.2 Thermodynamics of massive Maxwell-Boltzmann particles
$f_{eq}=\exp\left(-\frac{p_{0}-\mu_{0}}{T_{0}}\right)=\exp\left(-\frac{\sqrt{\vec{p}^{2}+m^{2}}-\mu_{0}}{T_{0}}\right).$ (9)
@@
$\varepsilon_{eq}=T_{eq}^{\alpha\beta}u_{\alpha}u_{\beta}=\int\,\frac{d^{3}p}{(2\pi)^{3}}\,\frac{(p\cdot u)^{2}}{p^{0}}f_{eq}=\frac{m^{2}T_{0}}{2\pi^{2}}\left(mK_{1}(\frac{m}{T_{0}})+3T_{0}K_{2}(\frac{m}{T_{0}})\right),$
$P_{eq}=-\frac{1}{3}\Delta_{\alpha\beta}T_{eq}^{\alpha\beta}=\frac{1}{3}\int\,\frac{d^{3}p}{(2\pi)^{3}}\,\frac{p^{2}}{p^{0}}f_{eq}=\frac{m^{2}T_{0}^{2}}{2\pi^{2}}K_{2}(\frac{m}{T_{0}}),$
$n_{eq}=u_{\mu}j_{eq}^{\mu}=\int\,\frac{d^{3}p}{(2\pi)^{3}}\,\frac{p\cdot u}{p^{0}}f_{eq}=\frac{m^{2}T_{0}}{2\pi^{2}}K_{2}(\frac{m}{T_{0}}),$
$S_{eq}=\frac{\varepsilon_{eq}+P_{eq}}{T_{0}}=\frac{m^{2}}{2\pi^{2}}\left(mK_{1}(\frac{m}{T_{0}})+4T_{0}K_{2}(\frac{m}{T_{0}})\right)=\frac{\partial P_{eq}(T)}{\partial T}\bigg{|}_{T=T_{0}},$ (11)
@@
$\Theta\equiv\frac{(T^{\mu}_{\mu})_{eq}}{T_{0}^{4}}=\frac{\varepsilon_{eq}-3P_{eq}}{T_{0}^{4}}=\frac{x^{3}K_{1}(x)}{2\pi^{2}},$
$c_{s}^{2}=\frac{\partial P_{eq}}{\partial\varepsilon_{eq}}=\frac{xK_{3}(x)}{x^{2}K_{0}(x)+5xK_{1}(x)+12K_{2}(x)}.$ (12)
where $x=m/T_{0}$. In Fig. 1, we present the trace anomaly (left panel) and the speed of sound (right panel) as functions of $x$. At $x=0$ (the massless limit), we find $\Theta=0$ and $c_{s}^{2}=1/3$. In contrast, as $x\to\infty$ (the heavy quasiparticle limit), both $c_{s}^{2}$ and $\Theta$ approach zero. This indicates that conformal symmetry is restored in both the massless and infinitely massive limits. However, the trace anomaly reaches its maximum value, $\Theta_{\text{Max}}\sim 0.058$, at $x=2.386$, demonstrating a significant departure from scale invariance at this point.
Figure 1: Left: the trace anomaly parameter $\Theta$ as given in Eq. (II.2), and right: the speed of sound $c_{s}^{2}$ as functions of $x=m/T_{0}$.
@@ III Charge Diffusion Correlators
$\chi(m,T_{0})\bigg{|}_{m\ll T_{0}}=\frac{T_{0}^{2}}{2\pi^{2}}\,\left(2-\frac{1}{2}\left(\frac{m}{T_{0}}\right)^{2}+\mathcal{O}\left(\left(\frac{m}{T_{0}}\right)^{4}\right)\right),$
$\chi(m,T_{0})\bigg{|}_{m\gg T_{0}}=\frac{T_{0}^{2}}{2\pi^{2}}\,\exp\left(-\frac{m}{T_{0}}\right)\left(\frac{15}{8}\left(\frac{m}{T_{0}}\right)^{\frac{1}{2}}\sqrt{\frac{\pi}{2}}+\mathcal{O}\left(\left(\frac{m}{T_{0}}\right)^{\frac{3}{2}}\right)\right).$ (19)
@@
where $v=\frac{1}{\sqrt{1+\frac{m^{2}}{p^{2}}}}$ and $k=|\vec{k}|$. This demonstrates that the massive case leads to different correlation functions compared to the massless case [27]. From definitions of correlation functions given in Eq. (II.1), we obtain the retarded current-current two-point function as follows
$G_{00}(\omega,\vec{k})=-\frac{\delta n}{\delta A_{0}}\bigg{|}_{\delta A=\delta g=0}=\,\frac{\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{1}(v,\omega,\vec{k})}{1-\frac{1}{2ik\tau_{eq}\chi}\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}},$
$G_{03}(\omega,\vec{k})=-\frac{\delta n}{\delta A^{3}}\bigg{|}_{\delta A=\delta g=0}=\frac{\omega}{k}\,\frac{\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{1}(v,\omega,\vec{k})}{1-\frac{1}{2ik\tau_{eq}\chi}\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}},$
$G_{33}(\omega,\vec{k})=-\frac{\delta j_{3}}{\delta A^{3}}\bigg{|}_{\delta A=\delta g=0}=\frac{G_{03}(\omega,\vec{k})}{\chi}\,\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{2}(v,\omega,\vec{k})+\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{3}(v,\omega,\vec{k}),$
$G_{11}(\omega,\vec{k})=G_{22}(\omega,\vec{k})=-\frac{\delta j_{1}}{\delta A^{1}}\bigg{|}_{\delta A=\delta g=0}=\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{4}(v,\omega,\vec{k}).$ (22)
@@
The unknown $\left(\mathcal{I}_{i}(v,\omega,\vec{k}),\,i=1,\cdots,4\right)$ functions are defined in the Appendix. B. There is no way to calculate these integrals directly. The only way is to use some approximations. We follow the expansion strategy in small and large $x=m/T_{0}$. To do this, we need to know the expansion in “$x$” for the following function
$\frac{1}{v}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}=L(\omega,\vec{k})+\frac{x^{2}}{2\tilde{p}^{2}}\left(L(\omega,\vec{k})-\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$\hskip 120.92421pt-\frac{x^{4}}{8\tilde{p}^{4}}\left(L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(k^{2}\tau_{eq}^{2}-(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\mathcal{O}(\frac{x^{6}}{\tilde{p}^{6}}),$
$\frac{1}{v}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}=\frac{2ik\tau_{eq}}{1-i\tau_{eq}\omega}.$ (23)
@@
where $\tilde{p}=\frac{p}{T_{0}}$ and $L(\omega,\vec{k})\equiv\ln{\left(\frac{\omega-k+\frac{i}{\tau_{eq}}}{\omega+k+\frac{i}{\tau_{eq}}}\right)}$. Also, it is necessary to know
$v\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}=L(\omega,\vec{k})-\frac{x^{2}}{2\tilde{p}^{2}}\left(L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$\hskip 113.81102pt+\frac{x^{4}}{8\tilde{p}^{4}}\left(3L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(3k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\mathcal{O}(\frac{x^{6}}{\tilde{p}^{6}}),$
$v\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}=0.$ (24)
@@
$\mathcal{S}^{(l)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}$
$=\chi(x)\bigg{(}L(\omega,\vec{k})+\frac{xK_{1}(x)}{2K_{2}(x)}\left(L(\omega,\vec{k})-\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$+\frac{x^{2}K_{0}(x)}{8K_{2}(x)}\left(L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(k^{2}\tau_{eq}^{2}-(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\cdots\bigg{)},$
$\mathcal{S}^{(h)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}=\frac{2ik\tau_{eq}\chi(x)}{1-i\tau_{eq}\omega}.$ (25)
The superscript $l(h)$ refers to the low (high) limits of $x$. Similarly, integrals of expressions in Eq. (III) can be performed
$\mathcal{T}^{(l)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,v\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}$
$=\chi(x)\bigg{(}L(\omega,\vec{k})-\frac{xK_{1}(x)}{2K_{2}(x)}\left(L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$-\frac{x^{2}K_{0}(x)}{8K_{2}(x)}\left(3L(\omega,\vec{k})+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(4k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\cdots\bigg{)},$
$\mathcal{T}^{(h)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,v\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}=0.$ (26)

### v2

@@ Abstract
Within the relaxation time approximation under a constant mass profile, we investigate the collective dynamics of a system of massive relativistic particles described by the Maxwell-Boltzmann equilibrium distribution. We analytically derive the two-point retarded correlation functions for both charge and energy-momentum tensor components at arbitrary momentum and frequency. We expand our results in the limits of very small and very large mass-to-temperature ratios ($m/T$). Similar to the massless case, we identify a critical threshold in ($k\tau$) below which the correlators permit physical solutions. This behavior arises from a logarithmic branch cut in the spectral function. At higher momenta, solutions emerge significantly below this cut, corresponding to non-hydrodynamic modes. Our analysis demonstrates that hydrodynamic poles dominate in the strong coupling regime, while the weak coupling regime features a logarithmic branch cut extending along $\omega=k$ and $\omega=-k$. Notably, in the sound channel, finite mass modifies the standard propagating sound mode, converting it into a purely imaginary mode. In contrast, the shear channel exhibits modes that asymptotically converge to their massless counterparts. Additionally, we compute the transport coefficients for shear and bulk viscosity, along with higher-order gradient corrections up to third order, expressed as perturbative expansions in both the small and large ($m/T$) regimes.
@@ I Introduction
In this paper, we investigate the Boltzmann equation within the Relaxation Time Approximation (RTA) for a system of particles with a constant mass profile. The chemical potential of the particles is neglected in the background limit. Employing a variational approach, we derive retarded two-point correlation functions for the charge and energy-momentum tensor components at arbitrary $(\omega,k)$. To explore collective behaviors, we examine the small-momentum limits of these functions. For the sake of analytical tractability, we expand expressions for both small and large values of the mass-to-temperature ratio, $x=m/T$. We obtain collective modes as well as the diffusion constant for extreme $x$ limits in the current-current correlation function. In the energy-momentum sector, we derive the fluctuations of the energy-momentum components and the complete spectrum of retarded correlation functions. Up to certain contact terms, the Ward identities are satisfied. Furthermore, we compute the shear and bulk transport coefficients as well as higher-order transports in derivatives for both small and large $x$ values. The hydrodynamic poles in the shear and sound channels are identified, revealing that in the sound channel, all modes become purely imaginary. Additionally, we observe the emergence of a branch cut between the $\omega=k$ and $\omega=-k$ in the solutions in the weak coupling regime ($\tau_{eq}T\to\infty$), while in the strong coupling regime ($\tau_{eq}T\to\infty$), hydrodynamic poles arises above a critical $(k\tau_{eq})$. This is the consequence of the logarithmic branch cut for small momenta.
@@ II.2 Thermodynamics of massive Maxwell-Boltzmann particles
$f_{eq}=\exp\left(-\frac{p_{0}-\mu_{0}}{T}\right)=\exp\left(-\frac{\sqrt{\vec{p}^{2}+m^{2}}-\mu}{T}\right).$ (9)
@@
$\varepsilon_{eq}=T_{eq}^{\alpha\beta}u_{\alpha}u_{\beta}=\int\,\frac{d^{3}p}{(2\pi)^{3}}\,\frac{(p\cdot u)^{2}}{p^{0}}f_{eq}=\frac{m^{2}T}{2\pi^{2}}\left(mK_{1}(\frac{m}{T})+3TK_{2}(\frac{m}{T})\right),$
$P_{eq}=-\frac{1}{3}\Delta_{\alpha\beta}T_{eq}^{\alpha\beta}=\frac{1}{3}\int\,\frac{d^{3}p}{(2\pi)^{3}}\,\frac{p^{2}}{p^{0}}f_{eq}=\frac{m^{2}T^{2}}{2\pi^{2}}K_{2}(\frac{m}{T}),$
$n_{eq}=u_{\mu}j_{eq}^{\mu}=\int\,\frac{d^{3}p}{(2\pi)^{3}}\,\frac{p\cdot u}{p^{0}}f_{eq}=\frac{m^{2}T}{2\pi^{2}}K_{2}(\frac{m}{T}),$
$S_{eq}=\frac{\varepsilon_{eq}+P_{eq}}{T}=\frac{m^{2}}{2\pi^{2}}\left(mK_{1}(\frac{m}{T})+4TK_{2}(\frac{m}{T})\right)=\frac{\partial P_{eq}(T)}{\partial T}\bigg{|}_{T},$ (11)
@@
$\Theta\equiv\frac{(T^{\mu}_{\mu})_{eq}}{T^{4}}=\frac{\varepsilon_{eq}-3P_{eq}}{T^{4}}=\frac{x^{3}K_{1}(x)}{2\pi^{2}},$
$c_{s}^{2}=\frac{\partial P_{eq}}{\partial\varepsilon_{eq}}=\frac{K_{3}(x)}{xK_{2}(x)+3K_{3}(x)}.$ (12)
In Fig. 1, we present the trace anomaly (left panel) and the speed of sound (right panel) as functions of $x$. At $x=0$ (the massless limit), we find $\Theta=0$ and $c_{s}^{2}=1/3$. In contrast, as $x\to\infty$ (the heavy quasiparticle limit), both $c_{s}^{2}$ and $\Theta$ approach zero. This indicates that conformal symmetry is restored in both the massless and infinitely massive limits. However, the trace anomaly reaches its maximum value, $\Theta_{\text{Max}}\sim 0.058$, at $x=2.386$, demonstrating a significant departure from scale invariance at this point.
Figure 1: Left: the trace anomaly parameter $\Theta$ as given in Eq. (II.2), and right: the speed of sound $c_{s}^{2}$ as functions of $x=m/T$.
@@ III Charge Diffusion Correlators
$\chi(m,T_{0})\bigg{|}_{m\ll T_{0}}=\frac{T_{0}^{2}}{2\pi^{2}}\,\left(2-\frac{1}{2}x^{2}+\mathcal{O}\left(x^{4}\right)\right),$
$\chi(m,T_{0})\bigg{|}_{m\gg T_{0}}=\frac{T_{0}^{2}}{2\pi^{2}}\,e^{-x}\left(x^{\frac{3}{2}}\sqrt{\frac{\pi}{2}}+\mathcal{O}\left(x^{\frac{1}{2}}\right)\right).$ (19)
@@
where $v=\frac{p}{\sqrt{p^{2}+m^{2}}}$ and $k=|\vec{k}|$. This demonstrates that the massive case leads to different correlation functions compared to the massless case [27]. From definitions of correlation functions given in Eq. (II.1), we obtain the retarded current-current two-point function as follows
$G_{00}(\omega,k)=-\frac{\delta n}{\delta A_{0}}\bigg{|}_{\delta A=0}=\,\frac{\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{1}(v,\omega,k)}{1-\frac{1}{2ik\tau_{eq}\chi}\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}},$
$G_{03}(\omega,k)=-\frac{\delta n}{\delta A^{3}}\bigg{|}_{\delta A=0}=\frac{\omega}{k}\,\frac{\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{1}(v,\omega,k)}{1-\frac{1}{2ik\tau_{eq}\chi}\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}},$
$G_{33}(\omega,k)=-\frac{\delta j_{3}}{\delta A^{3}}\bigg{|}_{\delta A=0}=\frac{G_{03}(\omega,\vec{k})}{\chi}\,\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{2}(v,\omega,k)+\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{3}(v,\omega,k),$
$G_{11}(\omega,k)=G_{22}(\omega,k)=-\frac{\delta j_{1}}{\delta A^{1}}\bigg{|}_{\delta A=0}=\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\mathcal{I}_{4}(v,\omega,k).$ (22)
@@
The unknown $\left(\mathcal{I}_{i}(v,\omega,k),\,i=1,\cdots,4\right)$ functions are defined in the Appendix. B. There is no way to calculate these integrals directly. The only way is to use some approximations. We follow the expansion strategy in small and large $x=m/T_{0}$. To do this, we need to know the expansion in “$x$” for the following function
$\frac{1}{v}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}=L(\omega,k)+\frac{x^{2}}{2\tilde{p}^{2}}\left(L(\omega,k)-\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$\hskip 120.92421pt-\frac{x^{4}}{8\tilde{p}^{4}}\left(L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(k^{2}\tau_{eq}^{2}-(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\mathcal{O}(\frac{x^{6}}{\tilde{p}^{6}}),$
$\frac{1}{v}\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}=\frac{2ik\tau_{eq}}{1-i\tau_{eq}\omega}-\frac{2i\tilde{p}^{2}k^{3}\tau_{eq}^{3}}{3x^{2}(1-i\tau_{eq}\omega)^{3}}+\frac{2i\tilde{p}^{4}k^{3}\tau_{eq}^{3}\left(3k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)}{15x^{4}(1-i\tau_{eq}\omega)^{5}}+\mathcal{O}(\frac{\tilde{p}^{6}}{x^{6}}),$ (23)
@@
where $\tilde{p}=\frac{p}{T_{0}}$ and $L(\omega,k)\equiv\ln{\left(\frac{\omega-k+\frac{i}{\tau_{eq}}}{\omega+k+\frac{i}{\tau_{eq}}}\right)}$. Also, it is necessary to know
$v\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}=L(\omega,k)-\frac{x^{2}}{2\tilde{p}^{2}}\left(L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$\hskip 113.81102pt+\frac{x^{4}}{8\tilde{p}^{4}}\left(3L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(3k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\mathcal{O}(\frac{x^{6}}{\tilde{p}^{6}}),$
$v\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}=\frac{2i\tilde{p}^{2}k\tau_{eq}}{x^{2}(1-i\tau_{eq}\omega)}-\frac{2i\tilde{p}^{4}k\tau_{eq}\left(k^{2}\tau_{eq}^{2}+3(1-i\tau_{eq}\omega)^{2}\right)}{3x^{4}(1-i\tau_{eq}\omega)^{3}}+\mathcal{O}(\frac{\tilde{p}^{6}}{x^{6}}).$ (24)
@@
$\mathcal{S}^{(s)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}$
$=\chi(x)\bigg{(}L(\omega,k)+\frac{xK_{1}(x)}{2K_{2}(x)}\left(L(\omega,k)-\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$+\frac{x^{2}K_{0}(x)}{8K_{2}(x)}\left(L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(k^{2}\tau_{eq}^{2}-(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\cdots\bigg{)},$
$\mathcal{S}^{(l)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,\frac{1}{v}\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}$
$=\chi(x)\bigg{(}\frac{2ik\tau_{eq}}{1-i\tau_{eq}\omega}-\frac{2ik^{3}\tau_{eq}^{3}K_{3}(x)}{xK_{2}(x)(1-i\tau_{eq}\omega)^{3}}+\frac{2ik^{3}\tau_{eq}^{3}\left(3k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)K_{4}(x)}{x^{2}K_{2}(x)\,(1-i\tau_{eq}\omega)^{5}}\bigg{)}.$ (25)
The superscript $s(l)$ refers to the small (large) limits of $x$. Similarly, integrals of expressions in Eq. (III) can be performed
$\mathcal{T}^{(s)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,v\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\ll 1}$
$=\chi(x)\bigg{(}L(\omega,k)-\frac{xK_{1}(x)}{2K_{2}(x)}\left(L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)}{(1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}}\right)$
$-\frac{x^{2}K_{0}(x)}{8K_{2}(x)}\left(3L(\omega,k)+\frac{2ik\tau_{eq}(1-i\tau_{eq}\omega)\left(4k^{2}\tau_{eq}^{2}+5(1-i\tau_{eq}\omega)^{2}\right)}{\left((1-i\tau_{eq}\omega)^{2}+k^{2}\tau_{eq}^{2}\right)^{2}}\right)+\cdots\bigg{)},$
$\mathcal{T}^{(l)}(x;\omega,\vec{k})\equiv\int\frac{dp}{2\pi^{2}T_{0}}\,p^{2}f_{eq}(p)\,v\,\ln{\left(\frac{\omega-k\,v+\frac{i}{\tau_{eq}}}{\omega+k\,v+\frac{i}{\tau_{eq}}}\right)}\bigg{|}_{x\gg 1}$
$=\chi(x)\bigg{(}\frac{6ik\tau_{eq}K_{3}(x)}{xK_{2}(x)(1-i\tau_{eq}\omega)}-\frac{10ik\tau_{eq}\left(k^{2}\tau_{eq}^{2}+3(1-i\tau_{eq}\omega)^{2}\right)K_{4}(x)}{x^{2}K_{2}(x)(1-i\tau_{eq}\omega)^{3}}\bigg{)}.$ (26)